OpenStudy (anonymous):

Show an equation and a solution for the problem. At noon, a train leaves Bridgetown heading east at 90 mph to Cartertown. Another train leaves Bridgetown at 1:00 p.m., also heading east at 100 mph. In how many hours after the second train left will they pass each other?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

90(x+1)+100x=time until they pass.(y) solve for x 90x+90+100x=y 120x+90=y 120x=-90 x=-.75 i did something wrong. Harkirat hopefully will have the right answer.

OpenStudy (anonymous):

By 1 pm the first train will be 90 mile ahead of the second train Since the difference between the speeds is 10 mph, the second train will make up 10 miles every hour and so will take 9 hours to make up the 90 miles and hence catch up and pass the first train So the second train will pass the first train after 9 hours

OpenStudy (anonymous):

do you need it in the form of formulas???

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

for an equation you might right \[1+90t=100t\] since distance = rate times time, and they obviously meet at the same time

OpenStudy (anonymous):

that is wrong.

OpenStudy (anonymous):

\[90+90t=100t\] is right i believe

OpenStudy (anonymous):

What is wrong

OpenStudy (anonymous):

what i wrote

OpenStudy (anonymous):

first trains distance is 90 + 90t, second trains distance is 100t and they meet at the same time so you have \[90+90t=100t\]

OpenStudy (anonymous):

let the no of hours after which the second train catches the first one be = x hours By this time the first train (which left earlier) would have been travelling for (x+1) hours So, in (x+1) hours the first train will cover distance = 90(x+1) miles and distance covered by the second train in x hours = 100x miles But since the second train catches and passes the first after x hours, this means that above distances will be equal at that point So, we get 90(x+1) = 100x 90x + 90 = 100x 90 = 100x-90x 90 = 10x 90/10 = x so x=9 Hence the second train will pass the first train after 9 hours

OpenStudy (anonymous):

My Medal !!!!??????

OpenStudy (anonymous):

help me with my other question